# OccluTrack: Rethinking Awareness of Occlusion for Enhancing Multiple Pedestrian Tracking

### v1

@@ III-A Overview
Input: A video sequence $V$; object detecor Det; Detection score threshold $\theta_{det}$; Pose-guided reID PG_ReID; Kalman Filter with Abnormal Motion Suppression KF_AMS; Occlusion-aware distance measurement ODM; Threshold for initilizing new tracklets $\theta_{new\_track}$
@@
             //W:warp matrix from t-1 to t
@@
III-B Abnormal Motion Suppression
Partial Occlusion in Kalman Filter
In the update step of the Kalman Filter, we find that the observation $z_{t}$ is incorrect during partial occlusion due to abnormal detections, i.e., suddenly becoming small of objects’ bounding boxes caused by partial occlusion. We assume that $z_{t}$ is corrupted by an error $e_{t}$, i.e., $z^{\prime}_{t}=z_{t}+e_{t}$, where $z^{\prime}_{t}$ is the corrupted observation. Hence, the updated state estimate in Eq.(2) becomes:
$\tilde{x}_{t|t}=\hat{x}_{t|t-1}+K_{t}(z_{t}+e_{t}-H\hat{x}_{t|t-1})=\hat{x}_{t|t}+K_{t}e_{t},$ (3)
where $\tilde{x}_{t|t}$ is the corrupted state estimate disturbed by the $K_{t}e_{t}$, and $x=[x,y,w,h,\dot{x},\dot{y},\dot{w},\dot{h}]^{T}$ representing the position of the bounding box $[x,y,w,h]^{T}$ and its moving speed $[\dot{x},\dot{y},\dot{w},\dot{h}]^{T}$. During partial occlusion, the error $e^{\prime}_{t}=K_{t}e_{t}=[e^{\prime}_{xt},e^{\prime}_{yt},e^{\prime}_{wt},e^{\prime}_{ht},e^{\prime}_{\dot{x}t},e^{\prime}_{\dot{y}t},e^{\prime}_{\dot{w}t},e^{\prime}_{\dot{h}t}]^{T}$ will dominate the update process. Subsequently, this error will accumulate during full occlusion without observations. In this period, the Kalman filter’s state update only relies on the predictions from Kalman Filter. Here, we denote the duration for the full occlusion as $\tau$. As Eq.(1) indicates, the state is predicted by $\hat{x}_{t|t-1}=A\hat{x}_{t-1|t-1}$, where $A$ is the $n\times n$ transition matrix for the state from $t-1$ to $t$. Formally, $A$ can be elaborated for the linear motion as
When the person re-appears after occlusion of $\tau$ from $t0$, with only the prediction of the Kalman Filter, the estimate can be predicted like :
$\begin{split}\tilde{x}_{t0+\tau}&=A^{\tau}\tilde{x}_{t0}=A^{\tau}(\hat{x}_{t0}+K_{t0}e_{t0})\\ &=\hat{x}_{t0+\tau}+A^{\tau}K_{t0}e_{t0}=\hat{x}_{t0+\tau}+A^{\tau}e^{\prime}_{t0}.\end{split}$ (5)
According to the transition matrix $A$ and error $e^{\prime}_{t}$, we can obtain $A^{\tau}$ and $e^{\prime}_{t0}$ as
$A^{\tau}=\begin{bmatrix}1&0&0&0&\tau&0&0&0\\ 0&1&0&0&0&\tau&0&0\\ 0&0&1&0&0&0&\tau&0\\ 0&0&0&1&0&0&0&\tau\\ 0&0&0&0&1&0&0&0\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&1\\ \end{bmatrix}_{8\times 8},$ (6)
$e^{\prime}_{t0}=[e^{\prime}_{x0},e^{\prime}_{y0},e^{\prime}_{w0},e^{\prime}_{h0},e^{\prime}_{\dot{x}0},e^{\prime}_{\dot{y}0},e^{\prime}_{\dot{w}0},e^{\prime}_{\dot{h}0}]^{T}.$ (7)
Hence, the accumulated error $A^{\tau}e^{\prime}_{t0}$ can be calculated by
$\small\vspace{5pt}A^{\tau}e^{\prime}_{t0}=\begin{bmatrix}e^{\prime}_{x0}+\tau e^{\prime}_{\dot{x}0}\\ e^{\prime}_{y0}+\tau e^{\prime}_{\dot{y}0}\\ e^{\prime}_{w0}+\tau e^{\prime}_{\dot{w}0}\\ e^{\prime}_{h0}+\tau e^{\prime}_{\dot{h}0}\\ e^{\prime}_{\dot{x}0}\\ e^{\prime}_{\dot{y}0}\\ e^{\prime}_{\dot{w}0}\\ e^{\prime}_{\dot{h}0}\\ \end{bmatrix}$ (8)
From Eq.(8), we can observe that the bounding box $[x,y,w,h]^{T}$ is shifted by a fixed error $[e^{\prime}_{x0},e^{\prime}_{y0},e^{\prime}_{w0},e^{\prime}_{h0}]^{T}$ and an accumulated error $[\tau e^{\prime}_{\dot{x}0},\tau e^{\prime}_{\dot{y}0},\tau e^{\prime}_{\dot{w}0},\tau e^{\prime}_{\dot{h}0}]^{T}$ with a factor of $\tau$.
@@ III-B Abnormal Motion Suppression
It is important to ensure that the observation error should be properly modeled and accounted for in Kalman Filter. To suppress abnormal motions during partial occlusion, we propose an abnormal motion suppression mechanism into Kalman Filter to detect and remove abnormal motions caused by partial occlusion. In particular, at Frame $t$, we first adopt a memory buffer to collect recent $N+1$ bounding boxes of a tracked pedestrian from $(t-N)$ to $t$, denoted as $\{B^{t}\}_{t=t-N}^{t}$ where $B^{t}=[x,y,w,h]^{T}$ is a bounding box. Here, we apply a speed filter to calculate the averaged speed of the previous $(N-1)$ bounding boxes:
$\bar{v}=\frac{1}{N-1}\sum_{n=t-N+1}^{t-1}(B^{n}-B^{n-1}),$ (9)
The current speed can be calculated by $v^{t}=B^{t}-B^{t-1}$, where $v=[v_{x},v_{y},v_{w},v_{h}]^{T}$. We define the speed of the center point and aspect ratio as $v_{c}=[v_{x},v_{y}]^{T}$ and $v_{a}=[v_{w},v_{h}]^{T}$, respectively. Then, we calculate the difference between the normalized current speed and the normalized average speed in the center point and aspect ratio, i.e., $d_{c}=\|v_{c}^{t}\|-\|\bar{v}_{c}\|$ and $d_{a}=\|v_{a}^{t}\|-\|\bar{v}_{a}\|$. Finally, we obtain the speed difference for frame $t$ by $d^{t}=[d_{c}^{T};d_{a}^{T}]^{T}=[d_{x},d_{y},d_{w},d_{h}]^{T}$.
With $d^{t}$, we can set a threshold $\theta_{v}$ to detect the abnormal center point moving and scale changing of a tracked person. Afterward, we suppress the Kalman gain in the update step, Eq.(2), to handle the detected abnormal motions caused by the observation error ($e_{t}$) during partial occlusion. Formally, we calculate the suppression coefficients $\{\alpha_{x},\alpha_{y},\alpha_{w},\alpha_{h}\}$ for each element of $d^{t}$ by:
$\alpha_{i}=\begin{cases}1,\text{ if }d^{t}_{i}\leq\theta_{v}\\ \alpha_{0},\text{ otherwise}\end{cases},$ (10)
@@
$\tilde{x}_{t|t}=\hat{x}_{t|t-1}+\alpha K_{t}(z_{t}-H\hat{x}_{t|t-1})+\alpha K_{t}e_{t}.$ (11)
With the suppression coefficient, the model suppresses the state estimate by $\alpha<1$ when abnormal motions occur ($e_{t}>0$), more trusting the predicted state $\hat{x}_{t|t-1}$, and keeps the state estimate by $\alpha=1$ in normal motions ($e_{t}=0$).
While addressing the challenge of error accumulation within the Kalman Filter, OC-SORT [22] concentrates on updating the Kalman Filter solely when pedestrian tracking can resume post occlusion. This approach places emphasis on incorporating velocity costs derived from previous observations into the association cost for updates. However, a notable drawback of this method is its tendency to overlook the impact of abnormal observations. By solely relying on observations, OC-SORT fails to account for the potential inaccuracies introduced by these abnormal observations, especially during instances of partial occlusion. This limitation becomes evident through the resulting significant disparities in velocities, leading to inaccurate association costs and, consequently, inaccurate associations.
@@ III-C Pose-guided re-ID
In multiple pedestrian tracking, re-ID modules in prior works normally treated the bounding boxes equally using global features regardless of obscured pedestrians or not. They overlooked the significance of local features that contain more identity information, particularly during partial occlusion. Some methods [43] employed Mask-RCNN [44] to extract local features, but they are coupled with detectors and challenging to optimize. In contrast, our approach introduces a pose-guided re-ID module to adaptively fuse local and global features according to the confidence of body parts, generating more informative person appearance representations.
@@
As depicted in Fig.3, we combine a real-time pose estimator with the re-ID model to obtain and fuse local body-part and global feature embeddings. Given a high-score bounding box $B^{t}_{h}$, we crop the persons from the input image as the input for the re-ID model. From the figure, the lower-stream ResNeSt [45] extracts the appearance feature $f\in\mathbb{R}^{H\times W\times C}$, where $H$ and $W$ denote the height and width of the feature map, and $C$ represents the channel size. Simultaneously, the upper-stream fast pose estimator [46] produces a heatmap $h\in\mathbb{R}^{H\times W\times 17}$ after Gaussian blur, indicating 17 keypoint positions and confidence scores. To reduce computation complexity, we group the heatmap $h\in\mathbb{R}^{H\times W\times 17}$ into the body-part heatmap $h^{\prime}\in\mathbb{R}^{H\times W\times 6}$ by summation, as illustrated in Fig.4. Consequently, we use the heatmap to create six masks $hm_{i}\in\mathbb{R}^{H\times W\times 1}$, where $i\in[0,5]$. Each mask is repeated in channels, producing a mask of $M_{i}\in\mathbb{R}^{H\times W\times C}$. Six body-part feature maps $f_{i}^{l}$ are obtained by $f_{i}^{l}=f\otimes M_{i}$, where $\otimes$ is the Hadamard product. The body-part feature embedding $E^{l}_{i}\in\mathbb{R}^{C}$ is generated by $E^{l}_{i}=\sum_{h=1}^{H}\sum_{w=1}^{W}f^{l}_{i}(h,w,c)$.
Figure 4: Combination of body keypoints to form different body parts. The keypoints are grouped into six body parts according to given rules.
We design local and global necks to extract the local and global feature embeddings. In the local neck, we concatenate the six body-part feature embeddings into $E^{l}_{c}=[(E^{l}_{0})^{T};(E^{l}_{1})^{T};...;(E^{l}_{5})^{T}]^{T}$ and employ a fully-connected layer to project the channels from $6\times C$ to $C$, resulting in a local feature embedding $E^{l}$. Conversely, the global neck applies the global average pooling for the appearance feature $f$, producing a global feature embedding $E^{g}\in\mathbb{R}^{C}$. During training, we employ two classifiers with batch normalization for the global and local feature embeddings. The Cross-Entropy loss and Triplet loss serve as the loss functions for each classifier, and we sum the four losses to form the overall loss during training:
@@ III-D Occlusion-aware Distance Measurement
In state-of-the-art trackers, like BoT-SORT [10], the IoU and appearance feature distances are performed conditionally. Specifically, the IoU and appearance embedding distances are first calculated between the high-score detections with feature embeddings and the tracklets. Conditioning on the IoU distance larger than a threshold, the final distance is chosen from the smaller one between the IoU distance and appearance feature distance. However, the IoU distance is usually inaccurate because of error accumulation during partial occlusion, as mentioned in Section III-B.
@@
To address the issue, we treat the observed tracklets and occluded tracklets in different manners. We adopt the occlusion-aware distance thresholds for two different types of tracklets. For observed tracklets, we keep the same settings as in BoT-SORT. For occluded tracklets, we set a higher IoU distance threshold ($\theta^{iou}_{i}$) and make it easier to keep track of occluded persons when the bounding boxes re-appear. Formally,
$\theta^{iou}_{i}=\begin{cases}\theta_{0}\text{ if }i\in I_{t}\\ \theta_{0}+o\text{ if }i\in I_{ut}\end{cases},$ (14)
where $o>0$ is the offset for the distance threshold, $i$ is the index for all tracklets, $I_{t}$ includes indices for tracked tracklets, $I_{ut}$ includes indices for untracked tracklets, and $\theta^{iou}_{i}$ represents the IoU distance threshold for Tracklet $i$. Then, we calculate the IoU distance $d^{iou}_{i,j}$ between Tracklet $i$ and Detection $j$ and the appearance embedding distance $\hat{d}^{cos}_{i,j}$ by:
@@ IV-A Dataset
We evaluate our proposed method on two multiple pedestrian tracking datasets, MOT17 and MOT20. The MOT17 dataset consists of 14 training and 12 testing video sequences, with over 7 hours of video footage. The videos were captured in various real-world scenarios, such as outdoor scenes, crowded public places, and indoor environments. The dataset contains more than 22,000 annotated pedestrian trajectories, with varying levels of occlusion, illumination changes, and other challenging factors. The MOT20 dataset contains 14 training and 11 testing video sequences, with over 9 hours of video footage. The videos were captured in similar scenarios as MOT17 but with improved annotation quality. The dataset contains over 60,000 annotated object trajectories, with challenging scenarios such as object interactions, fast motion, and occlusions.
@@ IV-B Evaluation Metrics
For both MOT17 and MOT20 datasets, we adopt CLEAR metrics introduced in [47]. The metrics include multiple object tracking accuracy (MOTA), IDF1 score, higher-order tracking accuracy (HOTA), false positive (FP), false negative (FN), ID switch (IDSw), association accuracy (AssA), and association recall (AssR). More importantly, MOTA evaluates the detection performance, IDF1 tests the performance of tracking, and HOTA is a more balanced evaluation metric of detection and tracking. FP and FN indicate the total number of false positives and false negatives. IDSw is the number of ID switches in the whole video. AssA and AssR evaluate the ability of association performance in terms of accuracy and recall.
IV-C Experimental Settings
We adopt YOLOX as the object detector following the settings in our baseline, where YOLOX is trained on multiple datasets [48, 49, 50, 51]. As for the parameters in our proposed method, we set the $\alpha_{0}=0.2$ in Eq.(10). And we set the buffer size of the bounding boxes the same as the buffer size for feature embeddings. And we adopt the FastPose in Alphapose [46] as the pose estimator to extract the heatmaps. We adopt ResNest [45] as the appearance feature extractor in pose-guided re-ID. And we follow the training strategies in our baseline for our pose-guided re-ID module. We train the re-ID model for 60 epochs based on the first half of the MOT17 and MOT20 training sets. And during training, the pose estimator is frozen and only provides the heatmaps. In the occlusion-aware distance measurement, we set the offset $o=0.2$ in Eq.(14). For other parameters, we keep the same settings in our baseline, where the $\theta^{cos}=0.25$ and $\theta^{iou}=0.5$. Before submitting results to MOTChanllenge, we also perform interpolation on the trajectories following our baseline [10].

### v2

@@ III-B Abnormal Motion Suppression Kalman Filter
Input: A video sequence $V$; object detecor Det; Detection score threshold $\theta_{det}$; Pose-guided reID PG_ReID; Kalman Filter with Abnormal Motion Suppression AMS_KF; Occlusion-aware distance measurement ODM; Threshold for initilizing new tracklets $\theta_{new\_track}$
@@
             //W:warp matrix from k-1 to k
@@
It is important to ensure that the observation error is properly modeled and accounted for in the Kalman Filter. To suppress abnormal motions during partial occlusion, we propose an abnormal motion suppression mechanism in Kalman Filter to detect and remove abnormal motions caused by partial occlusion. In particular, at Frame $k$, we first adopt a memory buffer to collect recent $N+1$ bounding boxes of a tracked pedestrian from $(k-N)$ to $k$, denoted as $\{B^{k}\}_{k=k-N}^{k}$ where $B^{k}=[x,y,w,h]^{\top}$ is a bounding box. Here, we apply a speed filter to calculate the averaged speed of the previous $(N-1)$ bounding boxes:
$\bar{v}=\frac{1}{N-1}\sum_{n=k-N+1}^{k-1}(B^{n}-B^{n-1}),$ (7)
The current speed can be calculated by $v^{k}=B^{k}-B^{k-1}$, where $v=[v_{x},v_{y},v_{w},v_{h}]^{\top}$. We define the speed of the center point and aspect ratio as $v_{c}=[v_{x},v_{y}]^{\top}$ and $v_{a}=[v_{w},v_{h}]^{\top}$, respectively. Then, we calculate the difference between the normalized current speed and the normalized average speed in the center point and aspect ratio, i.e., $d_{c}=\|v_{c}^{k}\|-\|\bar{v}_{c}\|$ and $d_{a}=\|v_{a}^{k}\|-\|\bar{v}_{a}\|$. Finally, we obtain the speed difference for frame $k$ by $d^{k}=[d_{c}^{\top};d_{a}^{\top}]^{\top}=[d_{x},d_{y},d_{w},d_{h}]^{\top}$.
With $d^{k}$, we can set a threshold $\theta_{v}$ to detect the abnormal center point moving and scale changing of a tracked person. Afterward, we suppress the Kalman gain in the update step, Eq. (2), to handle the detected abnormal motions caused by the observation error ($e_{k}$) during partial occlusion. Formally, we calculate the suppression coefficients $\{\alpha_{x},\alpha_{y},\alpha_{w},\alpha_{h}\}$ for each element of $d^{k}$ by:
$\alpha_{i}=\begin{cases}1,\text{ if }d^{k}_{i}\leq\theta_{v}\\ \alpha_{0},\text{ otherwise}\end{cases},$ (8)
@@
$\tilde{s}_{k|k}=\hat{s}_{k|k-1}+\alpha G_{k}(z_{k}-H\hat{s}_{k|k-1})+\alpha G_{k}e_{k}.$ (9)
With the suppression coefficient, the model suppresses the state estimate by $\alpha<1$ when abnormal motions occur ($e_{k}>0$), more trusting the predicted state $\hat{s}_{k|k-1}$, and keeps the state estimate by $\alpha=1$ in normal motions ($e_{k}=0$).
While addressing the challenge of error accumulation within the Kalman Filter, OC-SORT [25] concentrates on updating the Kalman Filter solely when pedestrians can still be tracked after occlusion. This approach emphasizes incorporating velocity costs derived from previous observations into the association cost for updates. However, a notable drawback of this method is its tendency to overlook the impact of abnormal observations. By solely relying on observations, OC-SORT fails to account for the potential inaccuracies introduced by these abnormal observations, especially during instances of partial occlusion. This limitation becomes evident through the resulting significant disparities in velocities, leading to inaccurate association costs and, consequently, inaccurate associations.
@@ III-C Pose-Guided Re-ID
Re-ID has been introduced to mitigate the occlusion issue within tracking-by-detection methods. Existing methods extract the full-body features to calculate similarities for Re-ID. However, the full-body features get noisy when partial occlusion happens. To address this issue, we introduce a pose-guide Re-ID module to extract features by body parts rather than full-body.
@@
In multiple pedestrian tracking, Re-ID modules in prior works normally treated the bounding boxes equally using global features regardless of whether the pedestrians are obscured or not. They overlooked the significance of local features that contain more identity information, particularly during partial occlusion. Some methods [48] employed Mask-RCNN [49] to extract local features, but they are coupled with detectors and challenging to optimize. In contrast, our approach introduces a pose-guided Re-ID module to adaptively fuse local and global features according to the confidence of body parts, generating more informative person appearance representations.
@@
As depicted in Fig. 3, we combine a real-time pose estimator with the Re-ID model to obtain and fuse local body-part and global feature embeddings. Given a high-score bounding box $B^{k}_{h}$, we crop the persons from the input image as the input for the Re-ID model. From the figure, the lower-stream ResNeSt [50] extracts the appearance feature $f\in\mathbb{R}^{H\times W\times C}$, where $H$ and $W$ denote the height and width of the feature map, and $C$ represents the channel size. Simultaneously, the upper-stream fast pose estimator [51] produces a heatmap $h\in\mathbb{R}^{H\times W\times 17}$ after Gaussian blur, indicating 17 keypoint positions and confidence scores. To reduce computation complexity, we group the heatmap $h\in\mathbb{R}^{H\times W\times 17}$ into the body-part heatmap $h^{\prime}\in\mathbb{R}^{H\times W\times 6}$ by summation, as illustrated in Fig. 4. Consequently, we use the heatmap to create six masks $hm_{i}\in\mathbb{R}^{H\times W\times 1}$, where $i\in[0,5]$. Each mask is repeated in channels, producing a mask of $M_{i}\in\mathbb{R}^{H\times W\times C}$. Six body-part feature maps $f_{i}^{l}$ are obtained by $f_{i}^{l}=f\otimes M_{i}$, where $\otimes$ is the Hadamard product. The body-part feature embedding $E^{l}_{i}\in\mathbb{R}^{C}$ is generated by $E^{l}_{i}=\sum_{h=1}^{H}\sum_{w=1}^{W}f^{l}_{i}(h,w)$.
Figure 4: Combination of body keypoints to form different body parts. The keypoints are grouped into six body parts according to the given rules.
We design local and global necks to extract the local and global feature embeddings. In the local neck, we concatenate the six body-part feature embeddings into $E^{l}_{cat}=[(E^{l}_{0})^{\top};(E^{l}_{1})^{\top};...;(E^{l}_{5})^{\top}]^{\top}$ and employ a fully-connected layer to project the channels from $6\times C$ to $C$, resulting in a local feature embedding $E^{l}$. Conversely, the global neck applies the global average pooling for the appearance feature $f$, producing a global feature embedding $E^{g}\in\mathbb{R}^{C}$. During training, we employ two classifiers with batch normalization for the global and local feature embeddings. The Cross-Entropy loss and Triplet loss serve as the loss functions for each classifier, and we sum the four losses to form the overall loss during training:
@@ III-D Occlusion-aware Distance Measurement
In data association, existing methods treat occluded and visible objects equally using the same association measurement. However, predictions for occluded objects may not be as accurate as those for visible objects, caused by factors like partial occlusion. Thus, in our occlusion-aware distance measurement, we loosen the requirements for associating objects that have been occluded.
@@
In state-of-the-art trackers, like BoT-SORT [13], the IoU and appearance feature distances are performed conditionally. Specifically, the IoU and appearance embedding distances are first calculated between the high-score detections with feature embeddings and the tracklets. Conditioning on the IoU distance larger than a threshold, the final distance is chosen from the smaller one between the IoU distance and the appearance feature distance. However, the IoU distance is usually inaccurate because of error accumulation during partial occlusion, as mentioned in Section III-B.
@@
To address the issue, we treat the observed tracklets and occluded tracklets in different manners. We adopt the occlusion-aware distance thresholds for two different types of tracklets. For observed tracklets, we keep the same settings as in BoT-SORT. For occluded tracklets, we set a higher IoU distance threshold ($\theta^{iou}_{i}$) and make it easier to keep track of occluded persons when the bounding boxes re-appear. In practice, we treat inactive tracklets as occluded tracklets as the tracklet will be set to inactive when it is occluded. Formally,
$\theta^{iou}_{i}=\begin{cases}\theta_{0}\text{ if }i\in I_{k}\\ \theta_{0}+o\text{ if }i\in I_{uk}\end{cases},$ (12)
where $o>0$ is the offset for the distance threshold, $i$ is the index for all tracklets, $I_{k}$ includes indices for tracked tracklets, $I_{uk}$ includes indices for untracked tracklets, and $\theta^{iou}_{i}$ represents the IoU distance threshold for Tracklet $i$. Then, we calculate the IoU distance $d^{iou}_{i,j}$ between Tracklet $i$ and Detection $j$ and the appearance embedding distance $\hat{d}^{cos}_{i,j}$ by:
@@ IV-A Dataset
We evaluate our proposed method on three tracking datasets: MOT17 [21], MOT20 [22], and DanceTrack [23]. The MOT17 [21] dataset consists of 7 training and 7 testing video sequences. The videos were captured in various real-world scenarios, such as outdoor scenes, crowded public places, and indoor environments. The dataset contains more than 22,000 annotated pedestrian trajectories, with varying levels of occlusion, illumination changes, and other challenging factors. The MOT20 [22] dataset contains 4 training and 4 testing video sequences. The videos were captured in similar scenarios as MOT17 but with a larger crowd density. The dataset contains over 60,000 annotated object trajectories, with challenging scenarios such as crowdedness. DanceTrack [23] is a benchmark designed to track multiple objects with uniform appearances and diverse motions. The dataset consists of 100 videos, with 40 videos for training, 25 for validation, and 35 for testing.
@@ IV-B Evaluation Metrics
For all three datasets, we adopt CLEAR metrics [52]. The metrics include multiple object tracking accuracy (MOTA), IDF1 score, higher-order tracking accuracy (HOTA), and ID switch (IDSw). More importantly, MOTA quantifies the overall tracking accuracy by accounting for both false positives (FP) and false negatives (FN). IDF1 evaluates the tracking capability to maintain object identities consistently over time. And HOTA is a more balanced evaluation metric of detection and tracking. HOTA extends MOTA on assessment by incorporating higher-order tracking errors, including fragmentation and merging. IDSw quantifies the number of identity switches occurring throughout the video sequence.
IV-C Experimental Details
We adopt YOLOX as the object detector following the settings in our baseline, where YOLOX is trained on multiple datasets [53, 54, 55, 56]. As for the parameters in our proposed method, we set the $\alpha_{0}=0.2$ in Eq. (8). And we set the buffer size of the bounding boxes the same as the buffer size for feature embeddings, which is 30. We adopt the FastPose in Alphapose [51] as the pose estimator to extract the heatmaps. We adopt ResNest [50] as the appearance feature extractor in pose-guided Re-ID. And we follow the training strategies in our baseline for our pose-guided Re-ID module. We train the Re-ID model for 60 epochs based on the first half of the MOT17, MOT20, and DanceTrack training sets, respectively. During training, the pose estimator is frozen and only provides the heatmaps. In the occlusion-aware distance measurement, we set the offset $o=0.2$ in Eq. (12). For other parameters, we keep the same settings in our baseline, where the $\theta^{cos}=0.25$ and $\theta^{iou}=0.5$. Before submitting results to MOTChallenge, we also perform interpolation on the trajectories following our baseline [13].